OpenStudy (anonymous):

Name the property of real numbers illustrated by the equation. -4(x + 3) = -4x - 12 A. Associative Property of Addition B. Commutative Property of Addition C. Associative Property of Multiplication D. Distributive Property

OpenStudy (skullpatrol):

$$\Huge -4(x + 3) = -4x - 12$$ $$\Huge a(b + c) =ab +ac$$

OpenStudy (anonymous):

A

OpenStudy (skullpatrol):

nope

OpenStudy (anonymous):

o_o

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

B :)

OpenStudy (skullpatrol):

nope

OpenStudy (anonymous):

it can be C, the last choice is distributive?

OpenStudy (anonymous):

can't**

OpenStudy (skullpatrol):

guessing in math does not work

OpenStudy (skullpatrol):

try to think what is going on in the question

OpenStudy (anonymous):

i know, im sure it was commutative

OpenStudy (skullpatrol):

$$\Huge ab=ba$$

OpenStudy (anonymous):

hmm ok

OpenStudy (anonymous):

Multiplication?

OpenStudy (skullpatrol):

-4 is being distributed as a multiplier, just like "a" in a(b+c)=ab+ac.

OpenStudy (anonymous):

C. Associative Property of multiplication

OpenStudy (skullpatrol):

nope

OpenStudy (anonymous):

Ooo it is Associative Property of Addition

OpenStudy (skullpatrol):

nope

OpenStudy (skullpatrol):

-4 is being distributed as a multiplier, just like "a" in a(b+c)=ab+ac.

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):